Find the concentration of a solution of acid that was made by adding 75mL of a 20% acid concentration to 50mL of a 40% acid concentration?
Answer by Theo(13342)

x = number of mL of the first solution.
y = number of mL of the second solution.

.2x = number of mL of acid in the first solution.
.4y = number of mL of acid in the second solution.

the total mL in your final solution is x + y.

the total mL of acid in your final solution is .2x + .4y.

you know that x is equal to 75mL and you know that y is equal to 50mL.

you can replace x and y with those figures in your formulas as shown below:

x + y becomes 75 + 50 which becomes 125mL.

.2x + .4y becomes .2*75 + .4*50 which becomes 15 + 20 which becomes 35mL.

you have 35mL of acid in your final solution out of a total of 125mL in your final solution.

the percentage of acid is therefore equal to 35/125 = 7/25 = .28 * 100 = 28%.